OpenStudy (misssmartiez):

Check: (4^4)^x = 4^32 X equals 28 because when you multiply the exponent bases, isn't it addition?

candycove (candycove):

http://openstudy.com/study#/groups/Mathematics

OpenStudy (misssmartiez):

Oh crap thats why it seemed different.

OpenStudy (misssmartiez):

x'D I always go to OS feedback when I was at math when my tab restarts. I keep forgetting to check it. Thank you Candy,

candycove (candycove):

You're welcome.

TheSmartOne (thesmartone):

psssttt hint: \(\Large (a^b)^c = a^{bc}\) and if \(\Large a^b = a^c\) then b = c